OpenStudy (anonymous):

What is the value of b^2 - 4ac for the following equation? x^2 + 5x + 4 = 0 A.) 9 B.) 16 C.) 25

OpenStudy (anonymous):

5^2 - 4*1*4= 25-16=9

OpenStudy (johnweldon1993):

Well the 'a' 'b' and 'c' represent the coefficients of the polynomial... So a = 1 b = 5 c = 4 So \[\large b^2 - 4ac\] becomes \[\large 5^2 - 4(1)(4)\] \[\large 25 - 16 = ?\]

OpenStudy (anonymous):

9..!

OpenStudy (johnweldon1993):

Indeed!

OpenStudy (anonymous):

What about this one? 2x^2 - 2x - 1 = 0

OpenStudy (anonymous):

-2^2 - 4*2*-1 = 4- (-8) or 4+8=12

OpenStudy (anonymous):

And this one? 5x^2 + 7x = 6

OpenStudy (anonymous):

do the same thing. a=5, b=7, c=-6 because you'd move the 6 over to make the equation =0. so you'd have 5x^2+7x-6=0